# COMMA: A Communicative Multimodal Multi-Agent Benchmark

### v1

@@ Appendix C Additional Statistics
Solver Expert Average Number of Mistakes (↓) Button Dog Wire Who LED Memory Keypad Password Color Maze Overall Human GPT4V 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 GPT4o 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 7.00 0.00 0.64 InternVL8b 2.50 6.00 0.75 5.00 7.50 12.00 9.00 2.50 0.00 0.50 4.14 GPT4V Human 3.00 2.33 2.33 0.33 1.67 1.33 0.00 0.00 0.00 3.00 1.45 GPT4o 1.67 3.67 0.00 0.33 2.00 3.67 1.33 0.00 4.00 1.67 1.83 InternVL8b 3.00 1.50 2.67 1.00 6.33 6.33 3.33 5.67 0.00 0.00 3.15 GPT4V GPT4V 2.30 2.40 0.10 0.60 5.50 1.30 5.40 1.40 0.00 4.70 2.37 QwenVL7b InternVL8b 5.50 4.20 5.00 13.40 0.00 6.90 10.80 0.00 0.00 0.00 4.58 InternVL GPT4o 3.70 3.80 0.00 7.60 8.40 6.70 16.00 1.60 0.00 0.00 4.78 GPT4o GPT4o 4.10 5.10 0.20 2.50 6.80 3.70 12.00 15.20 0.00 0.00 4.96 QwenVL7b GPT4o 3.00 3.30 6.55 12.10 9.60 7.00 7.90 0.00 0.00 1.40 5.10 InternVL26b InternVL26b 4.10 2.80 2.90 11.50 11.30 8.00 14.00 0.00 0.00 2.40 5.70 Random InternVL 4.17 3.21 2.80 3.48 9.92 14.08 14.68 2.04 0.00 3.96 5.86 InternVL8b QwenVL7b 3.30 2.40 14.80 13.50 10.30 8.80 16.30 0.00 0.00 0.00 6.94 QwenVL7b QwenVL7b 3.10 4.30 11.80 17.20 11.60 9.10 12.60 0.00 0.00 0.00 6.97 QwenVL2b QwenVL2b 2.50 10.60 13.30 15.20 1.90 16.00 16.20 0.00 0.00 1.00 7.67 InternVL8b InternVL8b 4.50 3.80 12.60 11.00 13.10 17.20 12.00 12.50 0.00 0.00 8.67
Table 2: Average number of mistakes the solver made for various puzzles. We average the mistakes over 10, 3, and 100 independent runs of each puzzle for the AI-AI, AI-Human, and random settings. The overall column is an average across all the puzzles.
@@
Solver Expert Average Conversation Length (↓) Button Dog Wire Who LED Memory Keypad Password Color Maze Overall Human GPT4V 2.00 2.00 2.67 3.00 3.67 7.67 6.67 6.00 20.00 2.67 5.30 GPT4o 2.67 2.00 3.00 6.00 1.33 8.33 10.50 3.00 20.00 2.67 5.79 InternVL8b 4.00 7.00 3.00 9.50 18.50 20.00 19.50 13.00 20.00 16.00 12.38 GPT4o Human 3.67 10.00 2.33 2.33 17.00 13.33 6.67 11.33 20.00 20.00 10.67 GPT4V 8.33 5.67 6.67 2.33 15.67 15.33 2.00 18.00 15.50 20.00 10.79 InternVL8b 5.50 4.00 5.67 3.33 14.00 17.00 7.67 11.00 20.00 20.00 10.93 GPT4V GPT4V 8.20 5.40 2.20 3.40 15.10 15.80 12.90 15.60 20.00 18.00 11.66 Random InternVL 5.17 4.21 3.80 4.48 12.96 20.00 19.88 20.00 20.00 19.79 12.98 GPT4o GPT4o 9.70 8.90 2.20 6.90 20.00 19.00 14.70 20.00 20.00 20.00 14.14 InternVL GPT4o 5.20 6.90 2.60 12.30 20.00 17.90 20.00 20.00 20.00 20.00 14.49 QwenVL2b QwenVL2b 3.50 11.60 14.30 16.20 4.60 20.00 18.20 20.00 20.00 20.00 14.84 InternVL26b InternVL26b 8.10 5.90 5.80 16.40 15.90 18.60 18.70 20.00 20.00 20.00 14.94 InternVL8b QwenVL7b 4.30 4.70 16.40 17.00 15.60 20.00 18.40 20.00 20.00 20.00 15.64 InternVL8b InternVL8b 5.50 6.30 14.70 13.20 20.00 18.80 20.00 20.00 20.00 19.90 15.84 QwenVL7b GPT4o 9.80 9.00 10.45 15.90 20.00 18.10 16.80 20.00 20.00 20.00 15.95 QwenVL7b InternVL8b 7.80 6.00 12.50 14.40 20.00 19.00 20.00 20.00 20.00 20.00 15.97 QwenVL7b QwenVL7b 8.90 12.10 14.70 18.20 17.30 20.00 18.60 20.00 20.00 20.00 16.98
Table 3: Average conversation length for various puzzles. We average the length of agent dialogue over 10, 3, and 100 independent runs of each puzzle for the AI-AI, AI-Human, and random settings. The overall column is an average dialogue length across all the puzzles.
@@
We report additional metrics recorded during evaluation such as Average Mistakes (Table 2) and Conversation Length (Table 3)

### v2

@@ Appendix C Additional Statistics
Model Average Success Rate % (↑) Wire Telehealth Who LED Memory Keypad Password Color Maze Atm Overall Human 100 ± 0.0 60 ± 15.5 90 ± 9.5 20 ± 12.7 50 ± 15.8 40 ± 15.5 80 ± 12.7 0 ± 0.0 80 ± 12.7 100 ± 0.0 65.14 ± 4.6 GPT-4o 98 ± 1.4 61 ± 4.9 72 ± 4.5 12 ± 3.2 22 ± 4.1 7 ± 2.5 2 ± 1.4 0 ± 0.0 4 ± 2.0 47 ± 5.0 32.50 ± 1.5 Gemini 85 ± 3.6 44 ± 5.0 35 ± 4.8 29 ± 4.5 1 ± 1.0 12 ± 3.2 4 ± 2.0 0 ± 0.0 7 ± 2.5 27 ± 4.4 24.40 ± 1.4 GPT-4V 77 ± 4.2 47 ± 5.0 39 ± 4.9 7 ± 2.5 0 ± 0.0 9 ± 2.9 0 ± 0.0 0 ± 0.0 5 ± 2.2 19 ± 3.9 20.30 ± 1.3 QwenVL 56 ± 5.0 40 ± 4.9 26 ± 4.4 12 ± 3.2 3 ± 1.7 6 ± 2.4 0 ± 0.0 0 ± 0.0 1 ± 1.0 0 ± 0.0 14.40 ± 1.1 LLaMA 3.2 64 ± 4.8 3 ± 1.7 27 ± 4.4 11 ± 3.1 11 ± 3.1 10 ± 3.0 0 ± 0.0 0 ± 0.0 2 ± 1.4 0 ± 0.0 12.80 ± 1.1 Random 57 ± 5.0 3 ± 1.7 44 ± 5.0 14 ± 3.5 0 ± 0.0 1 ± 1.0 0 ± 0.0 0 ± 0.0 3 ± 1.7 0 ± 0.0 12.20 ± 1.0 InternVL 61 ± 4.9 0 ± 0.0 28 ± 4.5 9 ± 2.9 1 ± 1.0 6 ± 2.4 1 ± 1.0 0 ± 0.0 0 ± 0.0 0 ± 0.0 9.72 ± 0.9 LLaVA 1.6 41 ± 4.9 1 ± 1.0 25 ± 4.3 16 ± 3.7 1 ± 1.0 4 ± 2.0 0 ± 0.0 0 ± 0.0 0 ± 0.0 0 ± 0.0 8.80 ± 0.9
Table 2: Average success rate of the solver made for 100 sampled initializations of various puzzles. The overall column is an average across all 1000 puzzle initializations. Bolded values indicate the highest average success rate in the column.
@@
Model Average Mistake Rate % (↓) Wire Telehealth Who LED Memory Keypad Password Color Maze Atm Overall Human 0.10 ± 0.1 1.70 ± 0.4 0.20 ± 0.1 2.00 ± 0.3 0.50 ± 0.2 2.00 ± 0.5 0.10 ± 0.1 3.00 ± 0.0 0.40 ± 0.2 0.00 ± 0.0 0.93 ± 0.1 GPT-4o 0.27 ± 0.1 1.60 ± 0.1 0.69 ± 0.1 0.98 ± 0.1 1.65 ± 0.1 2.74 ± 0.1 1.50 ± 0.1 2.79 ± 0.1 0.17 ± 0.0 1.66 ± 0.1 1.41 ± 0.0 GPT-4V 0.70 ± 0.1 1.56 ± 0.1 1.51 ± 0.1 1.35 ± 0.1 2.22 ± 0.1 2.53 ± 0.1 1.46 ± 0.1 2.79 ± 0.1 1.65 ± 0.1 1.44 ± 0.1 1.72 ± 0.0 Gemini 0.82 ± 0.1 1.90 ± 0.1 1.46 ± 0.1 0.46 ± 0.1 2.95 ± 0.0 1.43 ± 0.1 2.57 ± 0.1 3.00 ± 0.0 1.41 ± 0.1 1.95 ± 0.1 1.79± 0.0 QwenVL 1.43 ± 0.1 1.84 ± 0.1 2.36 ± 0.1 2.65 ± 0.1 2.77 ± 0.1 2.72 ± 0.1 0.77 ± 0.1 2.75 ± 0.1 0.95 ± 0.1 0.18 ± 0.1 1.84 ± 0.0 LLaVA 1.6 1.84 ± 0.1 2.75 ± 0.1 1.72 ± 0.1 2.37 ± 0.1 2.99 ± 0.0 2.36 ± 0.1 2.05 ± 0.1 0.59 ± 0.1 0.16 ± 0.0 2.00 ± 0.0 1.88 ± 0.0 LLaMA 3.2 1.47 ± 0.1 2.27 ± 0.1 2.16 ± 0.1 2.74 ± 0.1 2.79 ± 0.1 2.77 ± 0.1 2.09 ± 0.1 2.76 ± 0.1 0.27 ± 0.1 0.71 ± 0.1 2.00 ± 0.0 Random 1.70 ± 0.1 2.96 ± 0.0 2.02 ± 0.1 2.76 ± 0.1 3.00 ± 0.0 2.97 ± 0.0 0.94 ± 0.1 3.00 ± 0.0 1.96 ± 0.1 0.18 ± 0.1 2.15 ± 0.0 InternVL 1.52 ± 0.1 3.00 ± 0.0 2.35 ± 0.1 2.76 ± 0.1 2.99 ± 0.0 2.81 ± 0.1 2.28 ± 0.1 2.68 ± 0.1 0.36 ± 0.1 2.93 ± 0.0 2.41 ± 0.0
Table 3: Average mistake rate of the solver made for 100 sampled initializations of various puzzles. The overall column is an average across all 1000 puzzle initializations. Bolded values indicate the lowest average mistakes in the column.
Model Average Conversation Length % (↓) Wire Telehealth Who LED Memory Keypad Password Color Maze Atm Overall Human 2.10 ± 0.2 4.60 ± 0.6 4.00 ± 1.0 7.40 ± 0.9 9.40 ± 0.2 3.50 ± 0.3 7.00 ± 0.7 6.30 ± 0.6 2.60 ± 0.5 5.56 ± 0.6 4.95 ± 0.3 Gemini 2.0 1.23 ± 0.2 2.52 ± 0.2 3.55 ± 0.2 4.64 ± 0.1 4.09 ± 0.2 3.32 ± 0.2 2.93 ± 0.1 1.92 ± 0.2 4.45 ± 0.1 5.75 ± 0.2 3.44 ± 0.1 GPT-4o 1.55 ± 0.1 3.17 ± 0.2 2.88 ± 0.1 4.91 ± 0.0 9.24 ± 0.1 1.44 ± 0.1 4.45 ± 0.1 2.17 ± 0.1 4.95 ± 0.0 4.40 ± 0.3 3.92 ± 0.1 GPT-4V 2.31 ± 0.2 2.60 ± 0.3 3.76 ± 0.2 4.56 ± 0.1 7.55 ± 0.2 1.88 ± 0.2 4.56 ± 0.1 2.57 ± 0.1 4.71 ± 0.1 6.68 ± 0.3 4.12 ± 0.1 InternVL 0.36 ± 0.1 2.46 ± 0.1 1.73 ± 0.1 2.80 ± 0.1 2.53 ± 0.1 2.15 ± 0.1 3.12 ± 0.1 2.34 ± 0.2 4.89 ± 0.1 5.69 ± 0.1 3.05 ± 0.1 LLaMA 3.2 1.31 ± 0.1 4.84 ± 0.4 2.43 ± 0.1 3.16 ± 0.1 4.03 ± 0.2 1.63 ± 0.1 2.95 ± 0.2 2.25 ± 0.2 4.71 ± 0.1 9.54 ± 0.1 3.69 ± 0.1 LLaVA 1.6 2.23 ± 0.1 1.78 ± 0.3 3.15 ± 0.1 1.82 ± 0.2 3.71 ± 0.1 2.29 ± 0.2 3.79 ± 0.1 4.22 ± 0.2 5.00 ± 0.0 10.00 ± 0.0 3.80 ± 0.1 QwenVL 1.75 ± 0.1 1.64 ± 0.1 2.58 ± 0.1 3.84 ± 0.1 4.71 ± 0.3 2.04 ± 0.1 4.97 ± 0.0 3.64 ± 0.1 4.83 ± 0.1 10.00 ± 0.0 4.00 ± 0.1 Random 1.27 ± 0.1 2.27 ± 0.0 1.46 ± 0.1 2.76 ± 0.1 2.95 ± 0.1 2.88 ± 0.1 9.92 ± 0.0 2.58 ± 0.1 8.59 ± 0.2 10.00 ± 0.0 4.47 ± 0.1
Table 4: Average conversation length between the solver and expert made for 100 sampled initializations of various puzzles. The overall column is an average across all 1000 puzzle initializations. Bolded values indicate the lowest average conversation length in the column.
@@
We report additional metrics recorded during evaluation such as Average Success Rate (Table 2), Mistake Rate (Table 3), and Conversation Length (Table 4)